# DNA and Human Language: Epigenetic Memory and Redundancy in Linear Sequence

### v1

@@ Abstract
DNA has long been described as the “language of life”, but whether it possesses formal linguistic properties remains unresolved. Here, we present the first empirical evidence that DNA sequences exhibit core linguistic features—redundancy—through comprehensive analysis of genomic and epigenetic datasets. By mapping DNA sequences into a linguistic feature space, we demonstrate that fixed-length (41 bp) DNA segments encode information analogously to human language, with redundancy contributing to signal stability in aqueous environments. Moreover, we provide the first evidence of one-dimensional epigenetic memory, revealing how linear DNA sequences can store and maintain epigenetic marks like 6mA methylation, independent of higher-order chromatin structure[1]. Our approach also resolves persistent challenges in genomic data processing by applying a tailored linguistic mapping strategy that significantly improves data cleaning and analytical accuracy. Together, these findings establish a novel scientific paradigm that connects genomic information encoding with linguistic theory, laying the foundation for future development of large language models (LLMs) specifically designed for DNA.
@@ 1 Introduction
Early studies demonstrated that DNA-binding enzymes primarily recognize specific DNA sequences through interactions within the major groove of the double helix[2]. This process can be conceptually compared to how information is encoded and interpreted in human languages, where DNA-binding enzymes act as ‘readers’ of the sequence-specific instructions embedded in DNA. While this analogy aids understanding, it raises an intriguing question: what parallels, if any, exist between DNA and human languages, such as Chinese and English? Moreover, is it possible to draw a formal connection between the two systems of information encoding? For many years, Scientists have dreamed of linking the ‘language of life’—DNA—with human languages to deepen our understanding of DNA’s encoding and transmission mechanisms[3]. In 1994, Eugene Stanley observed that the frequency distribution of symbols in DNA sequences conforms to Zipf’s law[4], a statistical feature also found in natural language. This similarity sparked an interest in exploring the relationship between DNA and language. If we view organisms as carriers of language, their various components might correspond to words, phrases, and sentences in linguistic structures. This analogy suggests that natural language processing (NLP) methods could be applied to decode DNA. From an information theory perspective, redundancy enhances the accuracy of information transmission. Redundant information refers to "extra information added to the original data to ensure the correct reception of a message even in the presence of noise and other disturbances [5]." Redundancy in communication means that hearing one word helps us predict the next, ensuring that occasional missed or misheard words do not impede understanding. Redundant languages are more robust and adaptable. This raises the question: does short fixed-length DNA, as the ‘language of life’, also exhibit redundancy? While this remains unresolved, redundancy phenomena—such as the ‘many-to-one’ relationship between codons and amino acids, multiple gene copies, and genetic compensation mechanisms [6, 7]—contribute to the stability of life processes. With rapid advances in artificial intelligence, deep learning models can now automatically identify key information based on linguistic features, contextual information, and biological context—despite redundancy. Researchers are increasingly applying natural language algorithms to analyze DNA sequence information. For instance, when predicting DNA methylation on N6-adenine (6mA) methylation sites, algorithms such as LSTM, BiLSTM, and Transformer have significantly improved accuracy[8, 9, 10]. However, in some cases, model performance remains limited (e.g. below 90%)[11], often due to dataset quality and preprocessing issues. Unlike natural languages, DNA consists of only four symbols, which poses unique challenges in data cleaning—a pressing issue in bioinformatics that remains unresolved.
To understand the information storage mechanisms of DNA sequences and overcome the challenges associated with cleaning DNA sequence datasets, we address these issues simultaneously. If DNA datasets could be manually cleaned in the same way as Chinese or English sentences, this would suggest that the information storage mechanism of DNA could be deciphered by humans. In this study, we utilize 6mA DNA methylation data to represent DNA, and employ the pre-trained model BERT as a tool. By mapping DNA to the feature space of the Chinese language through an artificial dictionary, we simulate the “transition state” of the bases during DNA methylation. Our results demonstrate that using Chinese to simulate the base “transition state” in DNA methylation reaction is feasible. Then, rule-based extraction was used to automatically extract methylation information from the DNA sequences, and manual cleaning of the DNA sequence datasets was achieved. Moreover, we discovered that fixed-length DNA sequences can store the same information multiple times, exhibiting redundancy similar to that found in human natural language. The redundancy of methylation information in fixed-length DNA sequences enhances error tolerance and stabilizes signal transmission under environmental interference. In summary, by integrating DNA information processing with natural language parsing, we propose an innovative solution for DNA dataset cleaning, reveal information storage mechanisms in fixed-length DNA sequences.
@@ 10 Discussion and conclusion
This study takes DNA adenine methylation as an example to map DNA into the Chinese language space through dictionary mapping, achieving for the first time a conversion between DNA and human language. It was found that DNA, like human language, exhibits redundancy. Specifically, DNA possesses an embedded redundant information storage mechanism that helps maintain dynamic stability in liquid media, thereby supporting stable information transfer between DNA and DNA-binding proteins. Moreover, an effective cleaning of the DNA dataset was accomplished. In conclusion, we have proposed a highly innovative approach that addresses the challenges of DNA dataset cleaning and information storage by integrating natural language processing. This approach attempts to understand DNA encoding information from a "linguistic" perspective, even in the absence of known recognition sequences. The discovery of this novel encoding mechanism is expected to offer new insights into biological information systems.
@@
When redundancy occurs in a system, it will bring composability to the system, effectively increasing the variety of possible system states. It can be understood that enzymes will have multiple "embrace" forms combined with DNA double-stranded[15, 16], and this different "embrace" situation also implies slight differences in enzyme efficiency. On the contrary, the elimination of redundancy reduces composability and minimizes the number of system states, potentially leading to unintended off-target effects. This redundancy in DNA gives it defense, growth, scalability, adaptability, and fault tolerance, enabling it to cope with various situations in the environment. This provides the most reasonable explanation for chromatin’s highly dynamic changes and partial destruction during DNA replication, as well as the stability of epigenetic markers that cells are able to maintain throughout the cell cycle[17, 18, 19, 20].
@@
The latest reports suggest that memory formation results from DNA damage followed by repair in response to external environmental stimuli[21]. This process aligns with the discovery of redundant DNA information highlighted in this research, and may represent the essence of epigenetic inheritance and memory[22]. Traits governed by DNA methylation represent a form of inheritance that is susceptible to environmental influences, can be acquired and transmitted across generations, though it may be transient over successive lineages[23, 24]. At present, the mechanisms underlying acquisition and loss remain unclear. Evidence suggests that environment-induced epigenetic mutations and changes are non-random, with the environment driving alterations in specific genomic regions[25]. Jovasevic’s research group found that mice exposed to electric shock (an environmental stressor) activated TLR9 inflammatory signaling, which subsequently induced an immune process within the cell, leading to DNA damage and its repair. Notably, this process did not result in pathogenicity but instead modified the DNA sequence to form memory20. Their analysis further highlighted that unmethylated CpG DNA sequences can activate TLR9 signaling and may also result from DNA demethylation. Jovasevic’s research group research on mice aligns with the general process of epigenetic acquisition: environmental stimuli, DNA modification, sequence alterations, and the formation of long-term memory. Numerous studies have shown that environmental stimuli can induce changes in DNA methylation, resulting in hypermethylation (increased DNA methylation) or hypomethylation (DNA demethylation)[26, 27, 28, 29]. Research indicates that in nondividing cells, DNA can activate DNA demethylation processes to dynamically respond to external stimuli and establish new cellular states[30]. The redundant information features of DNA methylation identified on fixed-length DNA sequences in this study suggest that the interaction between the environment and epigenetics can be interpreted through the dynamics of DNA methylation gain and loss, as illustrated in the diagram below. The acquisition of
Fig. 5: The impact of the environment on 6mA DNA methylation.(Figure was created with Biorender.com)
@@
epigenetic traits,such as DNA methylation occurs through a series of events: ongoing environmental stimuli trigger immune responses, induce DNA damage and repair, and result in the introduction of new nucleobases. Persistent environmental exposure amplifies DNA methylation signals, leading to redundancy. As this redundancy increases, DNA methyltransferase becomes more likely to methylate the sequence, thereby safeguarding its integrity. The greater the redundancy, the more durable the memory, ensuring the stable inheritance of essential information. Consequently, the offspring exhibit a consistent genetic predisposition toward a specific personality trait. As for the epigenetic loss of features—characterized by DNA demethylation—this involves the loss of information and reduction of redundancy. Brief environmental stimuli often signify the dominance of another, more persistent environmental factor. In non-dividing cells, this process triggers repeated DNA demethylation, immune reactions, DNA damage, and repair, thereby simplifying the information encoded in DNA. It modifies individual nucleotides, disrupts pre-existing redundancy, reduces methylation signals, and complicates the labeling process by DNA methyltransferases, leading to off-target effects. Consequently, memory data is gradually erased, and traits governed by epigenetic mechanisms weaken across generations, while environmental pressures encode alternative trait information.
@@
Information transmission is inevitably subject to noise interference. To ensure accuracy, the sender must encode the information redundantly or repeat it, helping the receiver accurately interpret and complete the transmission. The discovery of redundancy mechanisms in DNA information marks only the beginning of deciphering genomic information. A vast amount of genomic information still needs to be explored and analyzed, which could deepen humanity’s understanding of itself and other life forms. In summary, our findings will provide significant insights and implications for fields such as genomics, synthetic biology, information storage, and gene editing.
@@ 4 Features description
In recent years, BERT has gained prominence as a pre-trained language representation model in natural language processing research. Its key feature is the unsupervised pre-training on large-scale corpora, which allows it to acquire linguistic knowledge embedded in the data. For instance, the Chinese pre-trained BERT model provides a feature vector space for Chinese characters, covering such information as characters, words, semantics, and syntax. Thus, we use the Chinese BERT model as a research tool to evaluate the performance of our Dataset 1.0 before and after cleaning by fine-tuning the model, and utilize deep learning to identify potential meaningful information in DNA sequences.

### v2

@@ Abstract
DNA is often described as the “language of life”, but whether it possesses formal linguistic properties remains unresolved. Here, we present the first empirical evidence that DNA sequences exhibit core linguistic features—specifically, functional and information redundancy—through comprehensive analysis of genomic and epigenetic datasets. By mapping DNA sequences into a linguistic feature space, we demonstrate that fixed-length (41 bp) DNA segments encode information analogously to human language, with redundancy contributing to signal stability in aqueous intracellular environments. Moreover, we provide the first evidence of one-dimensional epigenetic memory, showing that linear DNA sequences can maintain epigenetic marks such as 6mA methylation—contrasting with models focusing on epigenetic memory transmission via 3D chromatin organization[1]. Our tailored linguistic mapping strategy also addresses persistent challenges in genomic data processing, significantly improving data cleaning and feature extraction. Together, these findings establish a conceptual paradigm that bridges molecular information encoding and linguistic theory, laying the foundation for next-generation large language models (LLMs) specifically tailored to DNA—marking a shift at the interface of molecular biology, information theory, and artificial intelligence (AI).
@@ 1 Introduction
Early studies demonstrated that DNA-binding enzymes primarily recognize specific DNA sequences through interactions within the major groove of the double helix[2]. This mechanism can be conceptually compared to the encoding and decoding of information in human language, where DNA-binding enzymes act as “readers” that recognize and interpret specific instructions embedded within DNA sequences.While this analogy enhances our understanding of molecular recognition, it also raises a compelling question: Do DNA and human languages (such as Chinese or English) exhibit parallels in their informational structure or coding logic? More intriguingly, could a formal connection be established between these two systems of information encoding?
@@
For many years, Scientists have dreamed of linking the ‘language of life’—DNA—with human languages to deepen our understanding of DNA’s encoding and transmission mechanisms[3]. In 1994, Eugene Stanley observed that the frequency distribution of symbols in DNA sequences conforms to Zipf’s law[4], a statistical feature also found in natural language. This similarity sparked an interest in exploring the relationship between DNA and language. If we view organisms as carriers of language, their various components might correspond to words, phrases, and sentences in linguistic structures. This analogy suggests that natural language processing (NLP) methods could be applied to decode DNA. From an information theory perspective, redundancy enhances the accuracy of information transmission. Redundant information refers to "extra information added to the original data to ensure the correct reception of a message even in the presence of noise and other disturbances [5]." Redundancy in communication means that hearing one word helps us predict the next, ensuring that occasional missed or misheard words do not impede understanding. Redundant languages are more robust and adaptable. This raises the question: does short fixed-length DNA, as the ‘language of life’, also exhibit redundancy? While this remains unresolved, redundancy phenomena—such as the ‘many-to-one’ relationship between codons and amino acids, multiple gene copies, and genetic compensation mechanisms [6, 7]—contribute to the stability of life processes. With rapid advances in artificial intelligence, deep learning models can now automatically identify key information based on linguistic features, contextual information, and biological context—despite redundancy. Researchers are increasingly applying natural language algorithms to analyze DNA sequence information. For instance, when predicting DNA methylation on N6-adenine (6mA) methylation sites, algorithms such as LSTM, BiLSTM, and Transformer have significantly improved accuracy[8, 9, 10]. However, in some cases, model performance remains limited (e.g. below 90%)[11], often due to dataset quality and preprocessing issues. Unlike natural languages, DNA consists of only four symbols, which poses unique challenges in data cleaning—a pressing issue in bioinformatics that remains unresolved.
To understand the information storage mechanisms of DNA sequences and overcome the challenges associated with cleaning DNA sequence datasets, we address these issues simultaneously. If DNA datasets could be manually cleaned in the same way as Chinese or English sentences, this would suggest that the information storage mechanism of DNA could be deciphered by humans.In this study, we take 6mA methylation as an entry point to investigate the information-encoding logic embedded in DNA sequences. Using the BERT pre-trained model as a tool, we construct a custom dictionary to map DNA into the feature space of the Chinese language, thereby simulating the “transition state” of bases during the methylation process. Our results demonstrate that using Chinese to simulate the base “transition state” in DNA methylation reaction is feasible. The results indicate that short DNA segments contain functionally meaningful signal encodings with redundant features. This finding suggests that DNA harbors a symbolic information structure analogous to human language—capable of supporting memory maintenance, signal stability, and error self-correction within biological systems. It not only provides theoretical support for transcending traditional biochemical perspectives, but also lays the groundwork for constructing DNA-specific large language models that integrate symbolic encoding with molecular dynamics.
@@ 10 Discussion and conclusion
We mapped DNA adenine methylation into the linguistic space of Chinese, enabling the first semantic-level encoding of DNA into a human language. It was found that DNA, like human language, exhibits redundancy. Specifically, DNA possesses an embedded redundant information storage mechanism that helps maintain dynamic stability in liquid media, thereby supporting stable information transfer between DNA and DNA-binding proteins. Moreover, an effective cleaning of the DNA dataset was accomplished. In conclusion, we have proposed a highly innovative approach that addresses the challenges of DNA dataset cleaning and information storage by integrating natural language processing. This approach attempts to understand DNA encoding information from a "linguistic" perspective, even in the absence of known recognition sequences. The discovery of this novel encoding mechanism is expected to offer new insights into biological information systems.
To further synthesize our findings, we constructed a comparative framework highlighting the functional parallels between redundancy in human language and DNA sequences (Supplementary Table 27, Fig.5). In both systems, redundancy serves not merely as excess but as a functional mechanism enhancing error tolerance, adaptive flexibility, and context-dependent disambiguation. For instance, synonymous expressions in language ensure comprehension even under noisy conditions—mirrored by multiple redundant binding motifs in DNA that preserve gene regulation despite mutations. Likewise, linguistic substitutability (e.g., “I’m hungry” vs. “Shall we eat?”) parallels motif diversity in DNA, enabling dynamic regulatory responses under changing environments. These analogies suggest that redundancy is not only structurally embedded but also informationally optimized in both language and genome systems, contributing to stability, adaptability, and semantic nuance.
Fig. 5: Conceptual mapping of redundancy functions between human language and DNA. This model illustrates how three core types of redundancy in natural language—error tolerance, expressive flexibility, and contextual disambiguation—correspond to analogous mechanisms in DNA sequences. These include tolerance to mismatch, multiple motifs supporting the same biological function, and the regulatory role of sequence context. Together, these features contribute to a robust, redundant architecture that supports information stability, signal transmission, and epigenetic memory in biological systems.
@@
When redundancy occurs in a system, it will bring composability to the system, effectively increasing the variety of possible system states. It can be understood that enzymes will have multiple "embrace" forms combined with DNA double-stranded[15, 16], and this different "embrace" situation also implies slight differences in enzyme efficiency. On the contrary, the elimination of redundancy reduces composability and minimizes the number of system states, potentially leading to unintended off-target effects. This redundancy endows DNA with defense, scalability, and adaptability. It enables the system to cope with diverse and unpredictable environmental conditions. This provides the most reasonable explanation for chromatin’s highly dynamic changes and partial destruction during DNA replication, as well as the stability of epigenetic markers that cells are able to maintain throughout the cell cycle[1, 17, 18, 19].
@@
The latest reports suggest that memory formation results from DNA damage followed by repair in response to external environmental stimuli[20]. This process aligns with the discovery of redundant DNA information highlighted in this research, and may represent the essence of epigenetic inheritance and memory[21]. Traits governed by DNA methylation represent a form of inheritance that is susceptible to environmental influences, can be acquired and transmitted across generations, though it may be transient over successive lineages[22, 23]. At present, the mechanisms underlying acquisition and loss remain unclear. Evidence suggests that environment-induced epigenetic mutations and changes are non-random, with the environment driving alterations in specific genomic regions[24]. Jovasevic’s research group found that mice exposed to electric shock (an environmental stressor) activated TLR9 inflammatory signaling, which subsequently induced an immune process within the cell, leading to DNA damage and its repair. Notably, this process did not result in pathogenicity but instead modified the DNA sequence to form memory20. Their analysis further highlighted that unmethylated CpG DNA sequences can activate TLR9 signaling and may also result from DNA demethylation. Jovasevic’s research group research on mice aligns with the general process of epigenetic acquisition: environmental stimuli, DNA modification, sequence alterations, and the formation of long-term memory. Numerous studies have shown that environmental stimuli can induce changes in DNA methylation, resulting in hypermethylation (increased DNA methylation) or hypomethylation (DNA demethylation)[25, 26, 27, 28]. Research indicates that in nondividing cells, DNA can activate DNA demethylation processes to dynamically respond to external stimuli and establish new cellular states[29]. The redundant information features of DNA methylation identified on fixed-length DNA sequences in this study suggest that the interaction between the environment and epigenetics can be interpreted through the dynamics of DNA methylation gain and loss, as illustrated below(Fig. 6). The acquisition of
Fig. 6: Sequence-based information processing system for signal stability and cellular memory. This figure illustrates how environmental stimuli can affect DNA at the sequence level, including structural disruption, insertion or removal of redundant segments, and dynamic changes in epigenetic modifications such as 6mA methylation. Higher redundancy improves signal robustness under fluctuating conditions and supports the long-term maintenance of cellular memory, even in the presence of large-scale chromatin restructuring.(Figure was created with Biorender.com)
@@
epigenetic traits,such as DNA methylation occurs through a series of events: ongoing environmental stimuli trigger immune responses, induce DNA damage and repair, and result in the introduction of new nucleobases. Persistent environmental exposure amplifies DNA methylation signals, leading to redundancy. As redundancy increases, DNA methyltransferases are more likely to methylate the sequence, reinforcing its structural and functional integrity. In turn, higher redundancy leads to more durable memory, supporting the long-term inheritance of adaptive traits. Consequently, the offspring exhibit a consistent genetic predisposition toward a specific personality trait. As for the epigenetic loss of features—characterized by DNA demethylation—this involves the loss of information and reduction of redundancy. Brief environmental stimuli often signify the dominance of another, more persistent environmental factor. In non-dividing cells, this process triggers repeated DNA demethylation, immune reactions, DNA damage, and repair, thereby simplifying the information encoded in DNA. It modifies individual nucleotides, disrupts pre-existing redundancy, reduces methylation signals, and complicates the labeling process by DNA methyltransferases, leading to off-target effects. Consequently, memory data is gradually erased, and traits governed by epigenetic mechanisms weaken across generations, while environmental pressures encode alternative trait information.
@@
Information transmission is inevitably subject to noise interference. To ensure accuracy, the sender must encode the information redundantly or repeat it, helping the receiver accurately interpret and complete the transmission. The discovery of redundancy mechanisms in DNA information marks only the beginning of deciphering genomic information. A vast amount of genomic information still needs to be explored and analyzed, which could deepen humanity’s understanding of itself and other life forms. In summary, our findings will provide significant insights and implications for fields such as genomics, synthetic biology, information storage, and gene editing.More importantly, this framework offers a promising direction for developing cross-disciplinary tools that bridge molecular biology, natural language processing, and information theory—opening new avenues in precision medicine, genomic computing, and biological data engineering.Additionally, we observed that some redundant structures may exhibit potential functional directionality in other modification systems, suggesting that they might also play a role at the regulatory level.
@@ 4 Features description
In recent years, BERT has gained prominence as a pre-trained language representation model in natural language processing research. Its key feature is the unsupervised pre-training on large-scale corpora, which allows it to acquire linguistic knowledge embedded in the data. For instance, the Chinese pre-trained BERT model provides a feature vector space for Chinese characters, covering such information as characters, words, semantics, and syntax. Thus, we use thepre-trained Chinese BERT model (for character-level NLP) as a research tool to evaluate the performance of our Dataset 1.0 before and after cleaning by fine-tuning the model, and utilize deep learning to identify potential meaningful information in DNA sequences.